### v3

@@ 5.2 Data-driven reproduction number estimation strategy
The scale parameter, unknown in practice, is assumed constant in time. Remark that $\Phi_{t}(\boldsymbol{\mathsf{Z}})$ is linear in $\boldsymbol{\mathsf{Z}}$, with $\sum_{s\geq 1}\varphi_{s}=1$ by normalization of the serial interval distribution, hence if $s>t$, then $\partial_{\mathsf{Z}_{t}}\Phi_{s}(\boldsymbol{\mathsf{Z}})=\varphi_{s-t}$ is of order one. Thus, to ensure that $\alpha$ is reflecting the inflated variance observed in COVID-19 data while satisfying Assumption 2, the following data-driven heuristics is proposed $\alpha=0.1\times\mathrm{std}(\boldsymbol{\mathsf{Z}}).$ The regularization parameter $\lambda$ controlling the level of regularization in Equation (48) is selected in a data-driven manner through
@@ 5.3 Experimental setup
During the entire course of the COVID-19 pandemic, daily infection counts reported by National Health Agencies of 200+ countries worldwide were collected by Johns Hopkins University and made publicly available in an online repository.9 To demonstrate the universality of the proposed data-driven reproduction number estimation procedure, four countries, belonging to different continents, and two time periods, corresponding to late and early epidemic stages, are considered. Daily reported counts in France and India, from October 5, 2021 to February 6, 2023, and in Canada and Argentina, from December 22, 2020 to April 25, 2022, are downloaded from JHU repository. [^9] https://coronavirus.jhu.edu/ Each sequence of daily counts is then aggregated at the scale of the week, yielding a vector of observed counts $\boldsymbol{\mathsf{Z}}$ of length $T=70$ weeks. Finally, the associated global infectiousness $\boldsymbol{\Phi}(\boldsymbol{\mathsf{Z}})$ is computed using $\Phi_{t}(\boldsymbol{\mathsf{Z}})=\sum_{s\geq 1}\varphi_{s}\mathsf{Z}_{t-s}$, where $\varphi$ is the weekly discretized serial interval distribution introduced in Section 5.1. Top plots of Figure 4 shows weekly infection counts and associated global infectiousness in France 4(a) and India 4(b), from 2021 to 2023, and Canada 4(c) and Argentina 4(d), from 2020 to 2022.
@@
Figure 4: Data-driven estimation of the reproduction number of COVID-19 from weekly infection counts. Top rows: weekly aggregated infection counts. Middle rows: estimated effective reproduction number with data-driven hyperparameter selection based on the robustified $\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$. Bottom rows: robustified autoregressive unbiased prediction risk estimate averaged over $N=10$ realizations of the Monte Carlo vector with $95\%$ Gaussian confidence interval for a logarithmically spaced range of regularization parameters and minimizing regularization parameter obtained through grid search.
@@ 6 Conclusion and perspectives
A time-varying autoregressive model generalizing state-of-the-art viral epidemics models has been considered. The estimation of the parameters of the model, namely the time-varying reproduction coefficient, has been framed as a nonstandard, highly nonstationary, inverse problem. The proposed rigorous mathematical formulation enabled to leverage the efficient and versatile variational framework to design accurate reproduction coefficient estimators which are robust to high noise levels in the observations. The major contribution consisted in the design of asymptotically unbiased risk estimates, suited to the generalized autoregressive Poisson model, which are then plugged into an oracle strategy for fully data-driven fine-tuning of the hyperparameters of the variational estimator, removing the major obstacle to its practical use. The resulting data-driven time-varying estimation strategy has been assessed through intensive Monte Carlo simulations on synthetic data. Taking advantage of the proposed extended mathematical framework, the standard epidemiological model for viral epidemics has been enriched to account for the low quality of COVID-19 data, leading to a novel weekly scaled Poisson model. Finally, the data-driven time-varying estimation procedure is shown to yield very consistent estimation of the COVID-19 reproduction number in various countries and pandemic stages despite the low quality of reported data, demonstrating its practical applicability for epidemic monitoring in a crisis context. The data-driven nature of the proposed epidemiological indicator estimation strategy constitutes a major asset for its dissemination beyond signal processing as its use requires no expert knowledge.
Future work will consist in, first, extending the unbiased risk estimates to other noise models, e.g., to additive Gaussian and multiplicative Gamma noises inspiring from the generalized Stein estimators [23], in order to further enrich the proposed framework and to be able to extend the developed methodology to other models, in epidemiology or beyond. Second, as in epidemiology the memory functions are often parametric, with parameters encapsulating the pathogen transmission characteristics, the proposed framework will be leveraged to perform simultaneously the extraction of the memory functions parameters and the fine-tuning of the estimator hyperparameters.
For the sake of reproducibility and dissemination, a Matlab toolbox has been made publicly available on the GitHub of the corresponding author. [^10] https://github.com/bpascal-fr/APURE-Estim-Epi
7 Acknowledgements
B. Pascal is supported by ANR-23-CE48-0009 “OptiMoCSI” and S. Vaiter by ANR-18-CE40-0005 “GraVa”. The authors thank Patrice Abry for fruitful discussions.
@@ Appendix A Autoregressive Poisson Stein-like lemma
$\begin{split}\sum_{\mathsf{k}_{t}=0}^{\infty}\!\ldots\!\!&\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\prod_{s=1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}\\ &=\,\prod_{s=1}^{t-1}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}\times\sum_{k_{t}=0}^{\infty}\mathsf{G}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{t}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}\end{split}$ (51)
@@ Appendix B Autoregressive Poisson Unbiased Risk Estimate
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]$ =$\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left(\frac{\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\right)^{-t}\mathsf{Y}_{t}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{(\mathsf{Y}_{t}-\alpha_{t})\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\right].$ (68)

### v4

@@ 5.2 Data-driven reproduction number estimation strategy
The scale parameter, unknown in practice, is assumed constant in time. Remark that $\Phi_{t}(\boldsymbol{\mathsf{Z}})$ is linear in $\boldsymbol{\mathsf{Z}}$, with $\sum_{s\geq 1}\varphi_{s}=1$ by normalization of the serial interval distribution, hence if $s>t$, then $\partial_{\mathsf{Z}_{t}}\Phi_{s}(\boldsymbol{\mathsf{Z}})=\varphi_{s-t}$ is of order one. Thus, to ensure that $\alpha$ is reflecting the inflated variance observed in COVID-19 data while satisfying Assumption 2, the following data-driven heuristics is proposed $\alpha=0.1\times\mathrm{std}(\boldsymbol{\mathsf{Z}})$, where $\mathrm{std}(\boldsymbol{\mathsf{Z}})$ is the empirical standard deviation of $\boldsymbol{\mathsf{Z}}$. The regularization parameter $\lambda$ controlling the level of regularization in Equation (48) is selected in a data-driven manner through
@@ 5.3 Experimental setup
During the entire course of the COVID-19 pandemic, daily infection counts reported by National Health Agencies of 200+ countries worldwide were collected by Johns Hopkins University and made publicly available in an online repository.10 To demonstrate the universality of the proposed data-driven reproduction number estimation procedure, four countries, belonging to different continents, and two time periods, corresponding to late and early epidemic stages, are considered. Daily reported counts in France and India, from October 5, 2021 to February 6, 2023, and in Canada and Argentina, from December 22, 2020 to April 25, 2022, are downloaded from JHU repository. [^10] https://coronavirus.jhu.edu/ Each sequence of daily counts is then aggregated at the scale of the week, yielding a vector of observed counts $\boldsymbol{\mathsf{Z}}$ of length $T=70$ weeks. Finally, the associated global infectiousness $\boldsymbol{\Phi}(\boldsymbol{\mathsf{Z}})$ is computed using $\Phi_{t}(\boldsymbol{\mathsf{Z}})=\sum_{s\geq 1}\varphi_{s}\mathsf{Z}_{t-s}$, where $\varphi$ is the weekly discretized serial interval distribution introduced in Section 5.1. Top plots of Figure 4 show weekly infection counts and associated global infectiousness in France 4(a) and India 4(b), from 2021 to 2023, and Canada 4(c) and Argentina 4(d), from 2020 to 2022.
@@
Figure 4: Data-driven estimation of the reproduction number of COVID-19 from weekly infection counts. Top rows: weekly aggregated infection counts. Middle rows: estimated effective reproduction number with state-of-the-art EpiEstim [17, 74] in dark gray versus data-driven hyperparameter selection based on the robustified $\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ in red. Bottom rows: robustified autoregressive unbiased prediction risk estimate averaged over $N=10$ realizations of the Monte Carlo vector with $95\%$ Gaussian confidence interval for a logarithmically spaced range of regularization parameters and minimizing regularization parameter obtained through grid search.
@@ 6 Conclusion and perspectives
A time-varying autoregressive model generalizing state-of-the-art viral epidemics models has been considered. The estimation of the parameters of the model, namely the time-varying reproduction coefficient, has been framed as a nonstandard, highly nonstationary, inverse problem. The proposed rigorous mathematical formulation enables to leverage the efficient and versatile variational framework to design accurate reproduction coefficient estimators which are robust to high noise levels in the observations. The major contribution consists in the design of asymptotically unbiased risk estimates, suited to the generalized autoregressive Poisson model, which are then plugged into an oracle strategy for fully data-driven fine-tuning of the hyperparameters of the variational estimator, removing the major obstacle to its practical use. The resulting data-driven time-varying estimation strategy has been assessed through intensive Monte Carlo simulations on synthetic data. Taking advantage of the proposed extended mathematical framework, the standard epidemiological model for viral epidemics has been enriched to account for the low quality of COVID-19 data, leading to a novel weekly scaled Poisson model. Finally, the data-driven time-varying estimation procedure is shown to yield very consistent estimation of the COVID-19 reproduction number in various countries and pandemic stages despite the low quality of reported data, demonstrating its practical applicability for epidemic monitoring in a crisis context. The data-driven nature of the proposed epidemiological indicator estimation strategy and its low computational cost constitute major assets for its dissemination beyond signal processing as its use requires no expert knowledge.
Future work will consist in, first, implementing and assessing the unbiased risk estimates for the other noise models, namely additive Gaussian and multiplicative Gamma noises extending the generalized Stein estimators [23], in order to further enrich the proposed framework and to be able to apply the developed methodology to other models, in epidemiology or beyond. Second, as in epidemiology the memory functions are often parametric, with parameters encapsulating the pathogen transmission characteristics, the proposed framework will be leveraged to perform simultaneously the extraction of the memory functions parameters and the fine-tuning of the estimator hyperparameters.
For the sake of reproducibility and dissemination, a Matlab toolbox is publicly available on the GitHub of the corresponding author at https://github.com/bpascal-fr/APURE-Estim-Epi.
@@ Appendix A Autoregressive Poisson Stein-like lemma
$\sum_{\mathsf{k}_{t}=0}^{\infty}\!\ldots\!\!\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\prod_{s=1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}=\,\prod_{s=1}^{t-1}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}\times\sum_{k_{t}=0}^{\infty}\mathsf{G}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{t}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}$ (51)
@@ Appendix B Autoregressive Poisson Unbiased Risk Estimate
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]$ $\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left(\frac{\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\right)^{-t}\mathsf{Y}_{t}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{(\mathsf{Y}_{t}-\alpha_{t})\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\right].$ (68)